If the length of one side of a square is increased by 4 centimeters and the length of an adjacent side is multiplied by 4, the perimeter of the resulting rectangle is 3 times the perimeter of the square. Find the length of a side of the original square.
Answer by ankor@dixie-net.com(22740)

If the length of one side of a square is increased by 4 centimeters and the
length of an adjacent side is multiplied by 4, the perimeter of the resulting
rectangle is 3 times the perimeter of the square.
Find the length of a side of the original square.
:
Let x = the side of the square
then
(x+4) = one side of the rectangle
and
4x = the adjacent side of the rectangle
:
Rect perimeter = 3 time sq perimeter
2(x+4) + 2(4x) = 3(4x)
2x + 8 + 8x = 12x
10x + 8 = 12x
8 = 12x - 10x
8 = 2x
x = 4 cm is the side of the original square
:
:
Check this; (rect is 8 by 16)
2(8) + 2(16) = 3(16)
16 + 32 = 48